OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below: x • y • z = y • x • z Distributive Property Transitive Property Associative Property of Multiplication Commutative Property of Multiplication

OpenStudy (anonymous):

I believe it is the Commutative property of Multiplication

OpenStudy (anonymous):

Commutativity is the concept that \([a*b=b*a]\) or \([a+b=b+a]\)